Crew 280 Sol Summary Report 20-04-2023
Summary title: A Martian home in the Making
Author’s name: Mariona Badenas-Agustí, Commander.
Mission status: Nominal
Sol activity summary: As the Sun rose over the Martian desert, the aroma of freshly brewed coffee and warm, homemade bread filled the upper deck. Soon after, the crew gathered around the table to have a delicious breakfast and get ready for the day ahead.
During the morning, Neus Sabaté (Mission Specialist) managed to turn on her iron-based batteries in the Science Dome, while Cesca Cufí (Engineer) started building a container for her “Blob.” Meanwhile, Laia Ribas (GreenHab Officer) met with Núria Jar (Journalist) and Carla Conejo González (Executive Officer & Biologist) to discuss how to communicate her MDRS research with the public. In the observatory, Mariona Badenas-Agusti (Commander & Astronomer) aligned the telescope with Ariadna Farrés’ (Health & Safety Officer) help and was able to observe a beautiful Sunspot through the eyepiece (see Astronomy Report).
The crew reconvened for lunch, when they savored a fresh salad from the GreenHab and dehydrated scrambled eggs with vegetables. At 2PM, some crew members (Laia Ribas, Carla Conejo, Ariadna Farrés, and Mariona Badenas) started an EVA, heading to Cowboy Corner to fly a small drone, test our Midland radios, collect sand samples, and develop navigation strategies. The rest of the crew supported the EVA team and continued their research from the Hab. After returning to the base, everyone continued to work on their research.
As the Sun sets and SOL4 is drawing to a close, the crew is excited to prepare homemade pizzas with fresh tomatoes and basil from the GreenHab, complemented by a few dehydrated ingredients from the pantry. It has been a productive day, and Hypatia I looks forward to continuing their journey of exploration on the Red Planet. Building a home away from home may not be easy, but with a lot of work, effort, and (dough)-determination, anything is possible!
Look ahead plan: The crew will perform an EVA and continue their experiments throughout the day. Tomorrow, we will prepare some workshops for the weekend, including a celebration of Saint George’s special Catalan festivity.
Anomalies in work: Non-functional robotic telescope and Spirit needs a new battery wire.
Weather: Sunny, but high winds in the afternoon. Low 26F/High 55F.
Crew physical status: Nominal.
EVA: Not applicable today.
Reports to be filed: Sol Summary, EVA request, EVA Report, Astronomy Report, Operations Report, GreenHab Report, Journalist Report, Pictures of the Day.
Support Requested: (i) Replacement of stove knobs, if possible; (ii) retrieval of trash in the rear airlock.